OpenStudy (anonymous):

solve the equation : 3X^2 - 6X + 4 = 0

OpenStudy (anonymous):

no real solutions. are you using complex numbers?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

3(x^2 -2x) = -4 3(x^2 -2x +1) = -4 +3(1) (x-1)^2 = -1/3 x -1 = +- sqrt(-1/3) x = 1 +- sqrt(-1/3)

OpenStudy (amistre64):

\[x = 1 \pm i\frac{\sqrt{3}}{3}\] maybe